OpenStudy (anonymous):

AP Calculus: Find the point on the graph of the function that is closest to the given point. f(x)=(x+1)^2 Point (5,3)

OpenStudy (anonymous):

i dont know what to do

OpenStudy (jennychan12):

take the derivative of the function. then plug in 5

OpenStudy (anonymous):

find the normal to this curve and constrain it to pass the following point and point on the curve and this point dist is the answer

OpenStudy (anonymous):

ok the derivative of the function is: \[2(x+1) \times1\]

OpenStudy (anonymous):

@RajshikharGupta huh?

OpenStudy (jennychan12):

ok, then plug in 5

OpenStudy (anonymous):

\[2(5+1) \times 1\]

OpenStudy (jennychan12):

yep :)

OpenStudy (anonymous):

then what

OpenStudy (eyust707):

The function does not pass through that point. We want the line connecting our point points to be perpendicular to the slope f'(x).

OpenStudy (jennychan12):

you have a point and a slope.

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

how did u get m=10?

OpenStudy (jennychan12):

my bad sorry. point slope form (5,3) m =12

OpenStudy (anonymous):

how did u get m=12?

OpenStudy (anonymous):

i didnt knowwere supposed to use point slope form on these types of problems.

OpenStudy (jennychan12):

the derivative is the slope. your calc teacher should've told you that... u dont have to; it's just easier

OpenStudy (eyust707):

@jennychan12 that is not correct We have a function and some point which does not lie on the function  Which point on the function is closest to the other point in space?

OpenStudy (anonymous):

NO!NO! wait MAN what r u doing it is tangent not normal shortest distances r along normal

OpenStudy (eyust707):

I've never actually done this but I can' see it being that hard... I'll derive it if no one else does

OpenStudy (anonymous):

jennychan12 is wrong

OpenStudy (anonymous):

\[d=\sqrt{(x-5)^2 + ((x+1)^2-3)^2}\] \[d'=\frac{ 1 }{ 2 } ((x+5)^2 + (x^2 + 2x -2)^2)^\frac{ -1 }{ 2 } \times \left[ 2(x+5) \times 1 + 2(x^2 + 2x -2) \right]\]

OpenStudy (agent0smith):

One way to do this is using the distance formula. Then you find the minimum of that function, by differentiating it and finding where the derivative is equal to zero.

OpenStudy (anonymous):

thats what my teacher showed me ^

OpenStudy (anonymous):

just in place of 12 do-1/12 then evry thing after that is correct

OpenStudy (anonymous):

@agent0smith can u explain how he got that by using Distance formula? ^^"

OpenStudy (jennychan12):

sorry. i thought it was a tangent line problem.

OpenStudy (anonymous):

@agent0smith can u explain to me how he got that?

OpenStudy (agent0smith):

He's using the distance formula, and entering the points x2 - x1 and y2 - y1 into the formula. x1 = 5 and y1 = 3, and y2 you find in terms of x, by re-arranging your original f(x) equation. I'll write it out a little better, using the equation writer, so you can see.

OpenStudy (jennychan12):

just saying, no need to take derivative of the whole square root problem. just take the derivative of the inside. it'll give u the same answer. don't believe me? try it yourself.

OpenStudy (anonymous):

@jennychan12 yeah but I wanna try the method that my teacher tried to show me

OpenStudy (agent0smith):

Jenny is right, just minimize the d^2, no need for the square root. Since d^2 is a minimum when d is a minimum. distance formula is here: http://math.about.com/library/bldistance.htm \[x _{1}\] is just x, that's the point we want to find.\[x _{2}\] is 5. \[y _{2}\] is (x+1)^2 and \[y _{1}\] is 3.